OpenStudy (anonymous):

vector v has a horizontal vector component with magnitude 11 and a vertical component with magnitude 31 what is the acute angle formed by v and the positive x axis

OpenStudy (dumbcow):

 basically it makes a right triangle \[\tan \theta = \frac{31}{11}\] \[\theta = \tan^{-1} \frac{31}{11}\]